OpenStudy (anonymous):

If possible evaluate q+2r, q=[6, -1] r=[-3, 5] [-11, 8] [1, 0]

OpenStudy (anonymous):

2r + 5 = 19 Reorder the terms: 5 + 2r = 19 Solving 5 + 2r = 19 Solving for variable 'r'. Move all terms containing r to the left, all other terms to the right. Add '-5' to each side of the equation. 5 + -5 + 2r = 19 + -5 Combine like terms: 5 + -5 = 0 0 + 2r = 19 + -5 2r = 19 + -5 Combine like terms: 19 + -5 = 14 2r = 14 Divide each side by '2'. r = 7 Simplifying r = 7

OpenStudy (anonymous):

does this help